# A Topologically Enriched Probability Monad on the Cartesian Closed Category of CGWH Spaces

### v2

@@ 4. The image of $L$ under restriction of functions to $K_{\epsilon}$,
Let $X$ be a Polish space. Then following every Baire measure is a $k$-regular measure and hence, the notions of Baire, Radon and $k$-regular measure all coincide (see 4.3). Moreover, the topology on $\mathcal{P}(X)$ is given by the topology of weak convergence of measures and $\mathcal{P}(X)$ is again a Polish space.
@@
That every Baire measure is $k$-regular was already shown in Lemma 4.5. Now, let $\mathcal{P}_{w}(X)$ be the space of Baire (equivalently, Radon, by Theorem 4.3) measures on $X$ with the topology of weak convergence. As sets, $\mathcal{P}_{w}(X)=\mathcal{P}(X)$. We want to show that the respective topologies coincide, as well. First, note that the identity map $\mathcal{P}(X)\to\mathcal{P}_{w}(X)$ is continuous, since a convergent net in $\mathcal{P}(X)$ is also weakly convergent. By [BR07, p. 213, Theorem 8.9.4], our assumption that $X$ is a Polish space implies that $\mathcal{P}_{w}(X)$ is a Polish space, too. Hence, sequential continuity of the identity map $\mathcal{P}_{w}(X)\to\mathcal{P}(X)$, which is provided by Theorem 6.4, implies continuity, so the two topologies coincide and $\mathcal{P}(X)=\mathcal{P}_{w}(X)$ as topological spaces. ∎
@@ 6.4 The Baire probability monad is strongly affine
For the Baire probability monad $\mathcal{P}$ on $\mathsf{QCB}_{h}$, this will come down to the aforementioned phenomenon of deterministic marginal independence, as we now show.
@@
$\mathcal{B}a(X\times Y)=\mathcal{B}a(X)\otimes\mathcal{B}a(Y)=\sigma(\{A\times B\mid A\in\mathcal{B}a(X),B\in\mathcal{B}a(Y))\}).$

### v3

@@ 4. The image of $L$ under restriction of functions to $K_{\epsilon}$,
Let $X$ be a Polish space. Then every Baire measure is a $k$-regular measure and hence, the notions of Baire, Radon and $k$-regular measure all coincide (see 4.3). Moreover, the topology on $\mathcal{P}(X)$ is given by the topology of weak convergence of measures and $\mathcal{P}(X)$ is again a Polish space.
@@
That every Baire measure is $k$-regular was already shown in Lemma 4.5. Now, let $\mathcal{P}_{w}(X)$ be the space of Baire (equivalently, Radon, by Theorem 4.3) measures on $X$ with the topology of weak convergence. As sets, $\mathcal{P}_{w}(X)=\mathcal{P}(X)$. We want to show that the respective topologies coincide, as well. First, note that the identity map $\mathcal{P}(X)\to\mathcal{P}_{w}(X)$ is continuous, since a convergent net in $\mathcal{P}(X)$ is also weakly convergent. By [BR07, p. 213, Theorem 8.9.4], our assumption that $X$ is a Polish space implies that $\mathcal{P}_{w}(X)$ is a Polish space, too. Hence, sequential continuity of the identity map $\mathcal{P}_{w}(X)\to\mathcal{P}(X)$, which is provided by Theorem 6.4, implies continuity, so the two topologies coincide and $\mathcal{P}(X)=\mathcal{P}_{w}(X)$ also as topological spaces. ∎
@@ 6.4 The Baire probability monad is strongly affine
For the Baire probability monad $\mathcal{P}$ on $\mathsf{QCB}_{h}$, this will come down to the aforementioned phenomenon of deterministic marginal independence, as we now show. The proof relies heavily on the fact that product measures are well-behaved on QCB spaces (they agree with the usual measure-theoretic ones, see Lemma 5.7); this is not the case for general CGWH spaces.
@@
$\mathcal{B}a(X\times Y)=\mathcal{B}a(X)\otimes\mathcal{B}a(Y)=\sigma(\{A\times B\mid A\in\mathcal{B}a(X),B\in\mathcal{B}a(Y)\}).$